In an isentropic process, the volume of 500g of an ideal gas (cp = 0.52kJ/kg-K, cv= 0.32kJ/kg-K) was increased from 20L to 100L. If the initial pressure of the gas was 760kPa, compute for the final temperature in oK.

1
Expert's answer
2021-12-16T16:18:22-0500

"\gamma=\frac{c_p}{c_V}=1.6,"

"p_2=p_1(\frac{V_1}{V_2})^{\gamma}=55.6~kPa,"

"\nu=\frac{m(c_p-c_V)}R=12~mol,"

"T_2=\frac{p_2V_2}{\nu R}=55.6~K."